dx/ x² + a² = dy / xy -az = dz / xz + ay ( Using multipliers )

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
Using the multiplier theory we can write
          Dx/x^2+a^2 = dy/xy-az=dz/xz+ay= ldx+mdy+ndz/l(x^2+a^2)+m(xy-az)+ n(xz+ay)
         Letting (l,m,n)= (-y,x,a) solves the first part of the problem giving us,
       -ydx + xdy + adz=0 
Further integerating it with exact integerls:
we get,
xy + az = 0.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)